# Blowing up Dirac’s theorem

### v1

@@ Abstract.
We show that every graph $G$ on $n$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ is spanned by a complete blow-up of a cycle with clusters of quasi uniform size $\Omega(\log n)$. The proof is based on a recently introduced approach for finding vertex-spanning substructures via blow-up covers.
@@ 1. Introduction
A fundamental problem in extremal combinatorics is to determine the least number of edges that guarantees a copy of a graph $F$ in an $n$-vertex host graph $G$. Erdős and Stone [5] proved that host graphs whose density slightly exceeds the Turán density $\tau(F)$, meaning $e(G)\geq(\tau(F)+\varepsilon)\binom{n}{2}$, even contain a blow-up of $F$. Formally, such a blow-up is defined by replacing the vertices of $F$ with vertex-disjoint clusters and each edge with a complete bipartite graph whose parts are the corresponding clusters. The arguments of Erdős and Stone show that we can find blow-ups of $F$ with clusters of size $\Omega(\log n)$, which is optimal as witnessed by dense random graphs.
It is natural, then, to ask if this ‘blow-up’ phenomenon extends to other types of structures than copies of a fixed graph $F$. We study this question for vertex-spanning subgraphs. In this setting, one typically replaces conditions on the number of edges with conditions on the minimum degree. For instance, the classic theorem of Dirac [3] states that every graph $G$ on $n\geq 3$ vertices with $\delta(G)\geq n/2$ contains a Hamilton cycle. In analogy to the Erdős–Stone Theorem, Böttcher, Schacht and Taraz [1] proved that a minimum degree of $\delta(G)\geq(1/2+\varepsilon)n$ guarantees a vertex-spanning blow-up of a cycle with clusters of order $\Omega(1)$. We extend this with the following blown-up version of Dirac’s theorem.
@@
For every $\varepsilon,\eta>0$, there are $c>0$ and $n_{0}$ such that every graph $G$ on $n\geq n_{0}$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ contains a vertex-spanning blow-up of a cycle where each cluster has size in the range of $(1\pm\eta)c\log n$.
This result is related to recent work of the authors [7] for the analogous problem in hypergraphs and follows a similar proof strategy. However, unlike in the hypergraph setting, where the proof techniques limit us to clusters of $\operatorname{poly}(\log\log n)$ vertices, we are able to achieve asymptotically optimal cluster sizes in Theorem 1.1. It would be interesting to understand whether this can be extended to more complex structures such as powers of cycles.
@@ 2. Proof of the main result
Let us begin by introducing some terminology. Consider a set family $\mathcal{V}$, whose elements are assumed to be pairwise disjoint by convention of this paper. A set $X$ is $\mathcal{V}$-partite if it has at most one vertex in each part of $\mathcal{V}$. We say that $\mathcal{V}$ is $m$-balanced if $|V|=m$ for every $V\in\mathcal{V}$. Similarly, $\mathcal{V}$ is $(1\pm\eta)m$-balanced if $(1-\eta)m\leq|V|\leq(1+\eta)m$ for every $V\in\mathcal{V}$. Moreover, $\mathcal{V}$ is quasi $(1\pm\eta)m$-balanced if there is a singleton $V^{\ast}\in\mathcal{V}$ and $\mathcal{V}\smallsetminus\{V^{\ast}\}$ is $(1\pm\eta)m$-balanced. Given a graph $R$ and a set family $\mathcal{V}=\{V_{v}\}$ indexed by $v\in V(R)$, the $\mathcal{V}$-blow-up is the blow-up obtained by replacing each $v$ with the set $V_{v}$. We denote it by $R(\mathcal{V})$, and say that $R$ is the reduced graph of the blow-up. Some of the constant hierarchies are expressed in standard $\gg$-notation. To be precise, we write $y\gg x$ to mean that for any $y\in(0,1]$ there exists an $x_{0}\in(0,1)$ such that for all $x_{0}\geq x$ the subsequent statements hold. Hierarchies with more constants are defined in a similar way and are to be read from left to right following the order that the constants are chosen. Moreover, we tend to ignore rounding errors if the context allows it.
Our first lemma states that we can cover any Dirac-type graph with vertex-disjoint blow-ups whose reduced graphs inherit the minimum degree up to a small error.
@@
Let $\varepsilon,1/s\gg c,\eta\gg 1/n$ and $m=c\log n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Then the vertices of $G$ are covered by vertex-disjoint quasi $(1\pm\eta)m$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ has $s$ vertices and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/2)s$.
We note that Lemma 2.1 is a (quantitatively) refined version of a (qualitatively) more general result [7, Lemma 7.1]. The second lemma allows us to find a short chain of blow-ups, which connects two other clusters.
@@
The proofs of Lemmata 2.1 and 2.2 can be found in section 4 and section 6, respectively. We continue by deriving our main result.
@@
Introduce constants $s,c_{1},\eta,c_{2}$ with $\varepsilon\gg 1/s\gg c_{1},\eta\gg c_{2}\gg 1/n$ such that $\ell=c_{1}/c_{2}$ is an integer. (Note that the assumption $\varepsilon\gg\eta$ only gives a stronger outcome.) Set $m_{1}=c_{1}\log n$ and $m_{2}=c_{2}\log n$. Now consider an $n$-vertex graph $G$ with $\delta(G)\geq(1/2+\varepsilon)n$. Our goal is to find a $(1\pm 4\eta)m_{2}$-blow-up of a cycle, which allows us to conclude with $4\eta,c_{2}$ playing the rôles of $\eta,c$.
@@
By Lemma 2.1, the vertices of $G$ are covered by vertex-disjoint quasi-$(1\pm\eta)m_{1}$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ has order $s+1$ and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/2)(s+1)$. We shall find a spanning blow-up of a cycle in each of the blow-ups $R_{i}(\mathcal{V}_{i})$. Fix $1\leq i\leq t$, and let $V^{\ast}\in\mathcal{V}_{i}$ be the (unique) singleton cluster of $\mathcal{V}_{i}$ with $V^{\ast}=V_{v^{\ast}}$ for some $v^{\ast}\in V(R_{i})$. The graph $R^{\prime}_{i}=R_{i}-v^{\ast}$ has $s$ vertices, and $\delta(R^{\prime}_{i})\geq(1/2+\varepsilon/2)(s+1)-1\geq s/2$. By Dirac’s theorem, $R^{\prime}_{i}$ contains a Hamilton cycle $C^{\prime}_{i}$, say with vertices $c_{1},\dotsc,c_{s-1}$. Using again that $\delta(R^{\prime}_{i})\geq s/2$ together with the pigeonhole principle, we find $1\leq j<s$ such that both $c_{j}$ and $c_{j+2}$ are neighbours of $v^{\ast}$ (index computations modulo $s$). Let $\mathcal{V}^{\prime}_{i}$ be obtained from $\mathcal{V}_{i}$ by adding the vertex of $V^{\ast}$ to $V_{c_{j}}$ (keeping the name for convenience) and then removing $V^{\ast}$ from the family. Note then that $R^{\prime}_{i}(\mathcal{V}^{\prime}_{i})\subseteq G$ is a $(1\pm 2\eta)m_{1}$-balanced blow-up, where $R^{\prime}_{i}$ has $s$ vertices and a Hamilton cycle $C_{i}$. Let $\mathcal{V}^{\prime}$ be the union of the families $\mathcal{V}^{\prime}_{i}$ defined in this way over all $1\leq i\leq t$, and note that $\mathcal{V}^{\prime}$ is $(1\pm 2\eta)m_{1}$-balanced with $ts$ clusters.
Now we can build our spanning blown-up up cycle greedily, by merging two cycles at a time using Lemma 2.2. More precisely, we will construct such connections, in an iterative fashion, joining two blow-ups at a time using an $m_{2}$-balanced blow-up $P_{i}(\mathcal{K}_{i})$ where $P_{i}$ is path of order $3$. Given $1\leq i\leq t$, let us describe the $i$th step of the construction, where we join the blow-ups $\mathcal{V}^{\prime}_{i}$ and $\mathcal{V}^{\prime}_{i+1}$ respectively, with corresponding reduced graphs $R^{\prime}_{i},R^{\prime}_{i+1}$ (indices taken modulo $t$). We assume that in previous steps corresponding to $1\leq j<i$, we have already built connecting $m_{2}$-sized blow-ups with families $\mathcal{K}_{1},\dotsc,\mathcal{K}_{j}$ of three clusters each. Let $K=\bigcup_{1\leq j<i}\bigcup\mathcal{K}_{j}$ be the union of the vertices of all those clusters. Since $t\leq n/m_{1}$, we have that
$|K|\leq 3tm_{2}\leq 3\frac{n}{c_{1}\log n}c_{2}\log n=\frac{3}{\ell}n\leq\frac{\eta\varepsilon}{8}n.$
Suppose that $C_{i}=c_{1}^{i}\dotsb c_{s-1}^{i}$ and $C_{i+1}=c_{1}^{i+1}\dotsb c_{s-1}^{i+1}$ are two Hamilton cycles in $R^{\prime}_{i},R^{\prime}_{i+1}$, respectively. Let $V^{i}_{s-1}\in\mathcal{V}^{\prime}_{i}$ and $V^{i+1}_{1}\in\mathcal{V}^{\prime}_{i+1}$ be the clusters corresponding to vertices $c^{i}_{s-1}\in V(R^{\prime}_{i})$ and $c^{i+1}_{1}\in V(R^{\prime}_{2})$, respectively.
We define a set of vertices which we want to avoid when building $\mathcal{K}_{i}$. Let $U\subseteq V(G)$ contain the vertices of clusters $V\in\mathcal{V}^{\prime}$ with $|V\cap K|\geq\eta m_{1}$. Note that there are at most $|K|/(\eta m_{1})$ such clusters $V$. So
$|U|\leq(1+2\eta)m_{1}\frac{|K|}{\eta m_{1}}\leq 2\frac{|K|}{\eta}\leq\frac{\varepsilon}{4}n,$
which gives $|U\cup K|\leq\varepsilon n/4+\eta\varepsilon n/8\leq\varepsilon n/2$. This implies that $G\smallsetminus(U\cup K)$ has minimum degree at least $(1/2+\varepsilon/2)n$. Apply Lemma 2.2 with $(G\smallsetminus(U\cup K))\cup V^{1}_{s-1}\cup V^{2}_{1}$, $V^{1}_{s-1}$, $V^{2}_{1}$, $m_{1}$, $m_{2}$ playing the rôles of $G$, $U$, $V$, $m$, $m^{\prime}$ respectively. This yields $m_{2}$-sets $W_{1}\subseteq V^{1}_{s-1}$, $W_{2}\subseteq V^{2}_{1}$, and $W_{3}$, so that the complete bipartite graphs $K(W_{1},W_{3})$ and $K(W_{2},W_{3})$ are contained in $G$. We finish by setting $\mathcal{K}_{i}=\{W_{1},W_{2},W_{3}\}$ and let $P_{i}$ be the corresponding $3$-vertex path. By construction, at the end of this process we used at most $\eta m_{1}$ connecting vertices in each of the clusters of $\mathcal{V}^{\prime}$.
Set $\mathcal{K}=\bigcup_{1\leq i\leq t}\mathcal{K}_{i}$, and let $\mathcal{V}^{\prime\prime}$ be the set family obtained from $\mathcal{V}^{\prime}$ after removing the vertices of $\mathcal{K}$ from each cluster. By construction, $\mathcal{K}$ is $m_{2}$-balanced and $\mathcal{V}^{\prime\prime}$ is $(1\pm 3\eta)m_{1}$-balanced. Moreover, there is a cycle $C^{\prime}$ on $t(s-1)+t=ts$ vertices such that we can index the family $\mathcal{K}\cup\mathcal{V}^{\prime\prime}$ by $V(C^{\prime})$, so that $C^{\prime}(\mathcal{V}^{\prime\prime}\cup\mathcal{K})$ is a spanning blow-up of a cycle in $G$, with all clusters of size at least $m_{2}$.

### v2

@@ Abstract.
We show that every graph $G$ on $n$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ is spanned by a complete blow-up of a cycle with clusters of nearly uniform size $\Omega(\log n)$. The proof is based on a recently introduced approach for finding vertex-spanning substructures via blow-up covers.
Key words and phrases:
Hamilton cycles, minimum degree
2020 Mathematics Subject Classification:
05C35 (primary), 05C45 (secondary)
@@ 1. Introduction
A fundamental problem in extremal combinatorics is to determine the least number of edges that guarantees a copy of a graph $F$ in an $n$-vertex host graph $G$. Erdős and Stone [6] proved that host graphs whose density slightly exceeds the Turán density $\tau(F)$, meaning $e(G)\geq(\tau(F)+\varepsilon)\binom{n}{2}$, even contain a (complete) blow-up of $F$. Formally, such a blow-up is defined by replacing the vertices of $F$ with vertex-disjoint clusters and each edge with a complete bipartite graph whose parts are the corresponding clusters. The arguments of Erdős and Stone show that we can find blow-ups of $F$ with clusters of size $\Omega(\log n)$, which is optimal as witnessed by dense random graphs.
It is natural, then, to ask if this ‘blow-up’ phenomenon extends to types of structures other than copies of a fixed graph $F$. We study this question for vertex-spanning subgraphs. In this setting, one typically replaces conditions on the number of edges with conditions on the minimum degree. For instance, the classic theorem of Dirac [4] states that every graph $G$ on $n\geq 3$ vertices with $\delta(G)\geq n/2$ contains a Hamilton cycle. A consequence of the ‘Bandwidth Theorem’ of Böttcher, Schacht and Taraz [1] is that a minimum degree of $\delta(G)\geq(1/2+\varepsilon)n$ guarantees a vertex-spanning blow-up of a cycle with clusters of order $\Omega(1)$. We extend this with the following blown-up version of Dirac’s theorem.
@@
For every $\varepsilon,\eta>0$, there are $c>0$ and $n_{0}$ such that every graph $G$ on $n\geq n_{0}$ vertices with $\delta(G)\geq(1/2+\varepsilon)n$ contains a vertex-spanning blow-up of a cycle where each cluster has size in the range $(1\pm\eta)c\log n$.
This result is related to recent work of the authors [11] for the analogous problem in hypergraphs and follows a similar proof strategy. However, unlike in the hypergraph setting, where the proof techniques limit us to clusters of $\operatorname{poly}(\log\log n)$ vertices, we are able to achieve asymptotically optimal cluster sizes in Theorem 1.1.
Our result is related to the Bandwidth Theorem of Böttcher, Schacht and Taraz [1], which solved a conjecture of Bollobás and Komlós [8]. For the case of bipartite graphs, the theorem implies that a graph $G$ as in Theorem 1.1 contains every $n$-vertex bipartite graph $H$ that has ‘sublinear bandwidth’ and bounded maximum degree $\Delta=O(1)$. One can show that such a graph $H$ is contained in a blow-up of a path $B$ with clusters of size $o(n)$ (see [1, §7]), which explains the relation to embedding spanning graph blow-ups. Böttcher et al. [1] asked whether their result can be extended to graphs $H$ where $\Delta(H)$ grows with $n$. Partial progress in this question was made by Böttcher, Taraz and Würfl [2], who extended the result for graphs of ‘bounded arrangeability’ but not necessarily constant maximum degree. Theorem 1.1 answers this question without further assumptions on the maximum degree when the cluster sizes of $B$ are $o(\log n)$. It remains an open problem to determine whether this can be extended to sublinear cluster sizes (together with a suitable maximum degree condition).
Beyond this, it would be interesting to understand whether Theorem 1.1 can be generalised to more complex structures such as powers of cycles. Formally, the $r$th power of a cycle has a cyclically ordered set of vertices such that every $r$ consecutive vertices form an $r$-clique. It was shown by Komlós, Sárközy and Szemerédi [7] that, for all $k\geq 2$, the relative minimum degree threshold forcing a Hamilton $k$th power of a cycle is 11/(k+1)1-1/(k+1). We believe that our result extends as follows.
Conjecture 1.2.
For every $\varepsilon,\eta>0$ and $r\geq 1$, there are $c>0$ and $n_{0}$ such that every graph $G$ on $n\geq n_{0}$ vertices with $\delta(G)\geq(1-1/(r+1)+\varepsilon)n$ contains a vertex-spanning blow-up of the $r$th power of a cycle with cluster sizes in the range of $(1\pm\eta)c\log n$.
@@ 2. Proof of the main result
Let us begin by introducing some terminology. Consider a set family $\mathcal{V}$, whose elements are assumed to be pairwise disjoint by the convention of this paper. A set $X$ is $\mathcal{V}$-partite if it has at most one vertex in each part of $\mathcal{V}$. We say that $\mathcal{V}$ is $m$-balanced if $|V|=m$ for every $V\in\mathcal{V}$. Similarly, $\mathcal{V}$ is $(1\pm\eta)m$-balanced if $(1-\eta)m\leq|V|\leq(1+\eta)m$ for every $V\in\mathcal{V}$. Moreover, $\mathcal{V}$ is quasi $(1\pm\eta)m$-balanced if there is a singleton $V^{\ast}\in\mathcal{V}$ and $\mathcal{V}\smallsetminus\{V^{\ast}\}$ is $(1\pm\eta)m$-balanced. We refer to this unique singleton cluster as the exceptional cluster of $\mathcal{V}$. Given a graph $R$ and a set family $\mathcal{V}=\{V_{v}\}$ indexed by $v\in V(R)$, the $\mathcal{V}$-blow-up is the blow-up obtained by replacing each $v$ with the set $V_{v}$. We denote it by $R(\mathcal{V})$, and say that $R$ is the reduced graph of the blow-up. Our constant hierarchies are expressed in standard $\gg$-notation. To be precise, we write $y\gg x$ to mean that for any $y\in(0,1]$ there exists an $x_{0}\in(0,1)$ such that for all $x_{0}\geq x$ the subsequent statements hold. Hierarchies with more constants are defined in a similar way and are to be read from left to right following the order in which the constants are chosen. Moreover, we tend to ignore rounding errors if the context allows it.
Our main result is derived from the following two lemmata. The first one states that we can cover any Dirac-type graph with vertex-disjoint blow-ups whose reduced graphs inherit the minimum degree up to a minor deviation.
@@
Let $\varepsilon\gg 1/s\gg\eta\gg c\gg 1/n$ and $m=c\log n$. Let $G$ be an $n$-vertex graph with $\delta(G)\geq(1/2+\varepsilon)n$. Then the vertices of $G$ are covered by vertex-disjoint quasi $(1\pm\eta)m$-balanced blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ has $s$ vertices and satisfies $\delta(R_{i})\geq(1/2+\varepsilon/4)s$.
This result is reminiscent of the Regularity Lemma [14], which produces a single quasi-random blow-up with clusters of linear size and one cluster of exceptional vertices. In contrast, Lemma 2.1 gives a family of disjoint complete blow-ups with clusters of logarithmic size and one ‘exceptional’ singleton cluster. A quantitatively weaker result for hypergraphs can be found in our previous work [11, Lemma 7.1].
The second lemma for the proof of Theorem 1.1 allows us to find a short chain of blow-ups which connects two other clusters.
@@
The proofs of Lemmata 2.1 and 2.2 are based on a combination of subsampling with Erdős–Stone-type results and can be found in section 4 and section 6, respectively.
Let us continue with a proof overview on how to derive Theorem 1.1 from here. We begin by applying Lemma 2.1 to cover the vertex set of $G$ with disjoint blow-ups $R_{1}(\mathcal{V}_{1}),\dots,R_{t}(\mathcal{V}_{t})\subseteq G$ such that each reduced graph $R_{i}$ inherits the minimum degree condition and the clusters of each family $\mathcal{V}_{i}$ have size about $c_{1}\log n$ except for the exceptional cluster $V_{i}^{\ast}$, which is a singleton. Next, we carefully merge $V_{i}^{\ast}$ with another cluster while retaining the other properties and, for convenience, names. By Dirac’s theorem, each $R_{i}$ contains a Hamilton cycle $C_{i}$, which means that $R_{i}(\mathcal{V}_{i})$ is spanned by a blow-up of a cycle $C_{i}(\mathcal{V}_{i})$. In other words, we have found a partition of $V(G)$ into blow-ups of cycles $C_{1},\dots,C_{t}$ such that each cluster has size about $c_{1}\log n$. The idea is then to successively connect the cycles $C_{i}(\mathcal{V}_{i})$ and $C_{i+1}(\mathcal{V}_{i+1})$. More precisely, we apply Lemma 2.2 to find sets $\{U_{i}^{\prime},V_{i+1}^{\prime},W_{i}^{\prime}\}$ of size $c_{2}\log n$ each with $c_{1}\gg c_{2}$ such that the set $U_{i}^{\prime}$ is contained in a cluster of $C_{i}(\mathcal{V}_{i})$, the set $V_{i+1}^{\prime}$ is contained in a cluster of $C_{i+1}(\mathcal{V}_{i+1})$ and $G$ contains $K(U_{i}^{\prime},W_{i}^{\prime})$ and $K(V_{i}^{\prime},W_{i}^{\prime})$ as subgraphs. The sets $W_{i}$ may (arbitrarily) intersect with the rest of the graph, but we ensure that $\bigcup W_{i}$ only covers a small number of vertices of each cluster, which preserves balancedness. This results in a blow-up of a cycle that spans $G$, in which the ‘internal’ clusters have size approximately $c_{1}\log n$ and the ‘connecting’ clusters have size $c_{2}\log n$ as illustrated in Figure 1. We conclude by subdividing the internal clusters, winding around the reduced Hamilton cycle $C_{i}\subseteq R_{i}$, until all clusters have size about $c_{2}\log n$. Now come the details.
$V_{i}$$U_{i}$$V_{i+1}$$U_{i+1}$$R_{i}(\mathcal{V}_{i})$$R_{i+1}(\mathcal{V}_{i+1})$$U_{i}^{\prime}$$V_{i+1}^{\prime}$$W_{i}^{\prime}$$U_{i+1}^{\prime}$$W_{i+1}^{\prime}$$W_{i-1}^{\prime}$$V_{i}^{\prime}$
Figure 1. The sets of $\mathcal{K}_{i}=\{U_{i}^{\prime},V_{i+1}^{\prime},W_{i}^{\prime}\}$ connect the blow-ups $R_{i}(\mathcal{V}_{i})$ and $R_{i+1}(\mathcal{V}_{i+1})$.
@@
Introduce constants $s,c_{1},\eta,c_{2}$ with $\varepsilon\gg 1/s\gg c_{1},\eta\gg c_{2}\gg 1/n$ such that $\ell=c_{1}/c_{2}$ is an integer. (Note that the assumption $\varepsilon\gg\eta$ only gives a stronger outcome.) Set $m_{1}=c_{1}\log n$ and $m_{2}=c_{2}\log n$. Now consider an $n$-vertex graph $G$ with $\delta(G)\geq(1/2+\varepsilon)n$. Our goal is to find a $(1\pm 5\eta)m_{2}$-blow-up of a cycle, which allows us to conclude with $5\eta,c_{2}$ playing the rôles of $\eta,c$.
By Lemma 2.1, the vertices of $G$ are covered by vertex-disjoint quasi $(1\pm\eta)m_{1}$-balanced blow-ups $\hat{R}_{1}(\hat{\mathcal{V}}_{1}),\dots,\hat{R}_{t}(\hat{\mathcal{V}}_{t})\subseteq G$ such that each reduced graph $\hat{R}_{i}$ has order $s+1$ and satisfies $\delta(\hat{R}_{i})\geq(1/2+\varepsilon/4)(s+1)$. We shall find a spanning blow-up of a cycle in each of the blow-ups $\hat{R}_{i}(\hat{\mathcal{V}}_{i})$.
@@
To this end, we first merge the exceptional cluster with another cluster as follows. Fix $1\leq i\leq t$, and let $\hat{V}^{\ast}\in\hat{\mathcal{V}}_{i}$ be the (unique) singleton cluster of $\hat{\mathcal{V}}_{i}$ with $\hat{V}^{\ast}=\hat{V}_{v^{\ast}}$ for some $v^{\ast}\in V(\hat{R}_{i})$. The graph $R_{i}=\hat{R}_{i}-v^{\ast}$ has $s$ vertices, and $\delta(R_{i})\geq(1/2+\varepsilon/4)(s+1)-1>s/2$. By Dirac’s theorem, $R_{i}$ contains a Hamilton cycle $C_{i}$, say with vertices $v_{1},\dotsc,v_{s}$. Using again that $\delta(\hat{R}_{i})>s/2$ together with the pigeonhole principle, we find $1\leq j<s$ such that both $v_{j}$ and $v_{j+2}$ are neighbours of $v^{\ast}$ in $\hat{R}_{i}$ (index computations modulo $s$). Let $\mathcal{V}_{i}$ be obtained from $\hat{\mathcal{V}}_{i}$ by adding the vertex of $\hat{V}^{\ast}$ to the cluster of $v_{j+1}$ and then removing $\hat{V}^{\ast}$ from the family. Note then that $R_{i}(\mathcal{V}_{i})\subseteq G$ is a $(1\pm 2\eta)m_{1}$-balanced blow-up, where $R_{i}$ has $s$ vertices and a Hamilton cycle $C_{i}$. Let $\mathcal{V}$ be the union of the families $\mathcal{V}_{i}$ defined in this way over all $1\leq i\leq t$, and note that $\mathcal{V}$ is $(1\pm 2\eta)m_{1}$-balanced with $ts$ clusters.
Now we can build our spanning blown-up cycle greedily, by connecting two cycles at a time using Lemma 2.2. In each blow-up $\mathcal{V}_{i}$, fix two distinct clusters $U_{i}$ and $V_{i}$ whose corresponding vertices in $R_{i}$ are consecutive in $C_{i}$. The following claim allows us to connect each $R_{i}(\mathcal{V}_{i})$ with $R_{i+1}(\mathcal{V}_{i+1})$ (index computations modulo $t$).
Claim 2.3.
For $1\leq i\leq t$, there are $m_{2}$-balanced families $\mathcal{K}_{i}=\{U_{i}^{\prime},V_{i+1}^{\prime},W_{i}^{\prime}\}$ with $U_{i}^{\prime}\subseteq U_{i}$ and $V_{i+1}^{\prime}\subseteq V_{i+1}$ such that the complete bipartite graphs $K(U_{i}^{\prime},W_{i}^{\prime})$ and $K(V_{i+1}^{\prime},W_{i}^{\prime})$ are contained in $G$. In addition, setting $K_{i}=\bigcup_{1\leq j<i}\bigcup\mathcal{K}_{j}$ for each $1\leq i\leq t$, the following hold:
@@
    $K_{i}$ does not intersect any of the clusters $U_{j}$ for $i\leq j\leq t$ and $V_{j}$ for $i<j\leq t+1$ (index computations modulo $t$),
[…]
    $K_{i}$ intersects any cluster in at most $2\eta m_{1}$ vertices.
Proof of the claim.
For $1\leq i<t$, let us assume that we have already built connecting $m_{2}$-sized blow-ups with families $\mathcal{K}_{1},\dotsc,\mathcal{K}_{i}$ satisfying the claim. Since $t\leq n/(1-\eta)m_{1}\leq 2n/m_{1}$, we have that
$|K_{i}|\leq 3tm_{2}\leq 6\frac{n}{c_{1}\log n}c_{2}\log n=\frac{6}{\ell}n\leq\frac{\eta\varepsilon}{16}n\,.$
We define two sets of vertices which we want to avoid when building $\mathcal{K}_{i}$. Let $Y$ contain the vertices of the clusters $U_{j}$ for $i+1\leq j\leq t$ and $V_{j}$ for $i+1<j\leq t+1$. Note that $|Y|\leq 2n/((1-\eta)s)\leq\varepsilon n/8$, since $U_{i}$ and $V_{i}$ contain a share of at most $2/((1-\eta)s)$ vertices of $\mathcal{V}_{i}$. We let $X\subseteq V(G)$ contain the vertices of saturated clusters $V\in\mathcal{V}$, that is, those clusters $V$ such that $|V\cap K_{i}|\geq\eta m_{1}$. Note that there are at most $|K_{i}|/(\eta m_{1})$ saturated clusters. So
$|X|\leq(1+2\eta)m_{1}\frac{|K_{i}|}{\eta m_{1}}\leq 2\frac{|K_{i}|}{\eta}\leq\frac{\varepsilon}{8}n\,$,
which gives $|X\cup Y\cup K_{i}|\leq\varepsilon n/2$. This implies that $G^{\prime}=G-(X\cup Y\cup K_{i})$ has minimum degree at least $(1/2+\varepsilon/2)n$. Apply Lemma 2.2 with $G^{\prime}$, $U_{i}$, $V_{i+1}$, $m_{1}$, $m_{2}$ playing the rôles of $G$, $U$, $V$, $m$, $m^{\prime}$ respectively. This yields $m_{2}$-sets $U_{i}^{\prime}\subseteq V_{i}$, $V_{i+1}^{\prime}\subseteq V_{i+1}$, and $W_{i}^{\prime}$, so that the complete bipartite graphs $K(U_{i}^{\prime},W_{i}^{\prime})$ and $K(V_{i+1}^{\prime},W_{i}^{\prime})$ are contained in $G$. We finish by setting $\mathcal{K}_{i}=\{U_{i}^{\prime},V_{i+1}^{\prime},W_{i}^{\prime}\}$. Note that $K_{i+1}=K_{i}\cup\bigcup\mathcal{K}_{i}$ satisfies i. Moreover, $K_{i+1}$ does not intersect with saturated clusters at all. For any non-saturated cluster, $K_{i+1}$ intersects it in at most $2\eta m_{1}$ vertices, where $\eta m_{1}$ vertices account for the fact that the cluster is not saturated, and $3m_{2}\leq\eta m_{1}$ account for the vertices in $K_{i}$. This gives ii. ∎
Set $\mathcal{K}=\bigcup_{1\leq i\leq t}\mathcal{K}_{i}$, and let $\mathcal{V}^{\prime}$ be the set family obtained from $\mathcal{V}$ after removing the vertices of $\mathcal{K}$ from each cluster. By construction, $\mathcal{K}$ is $m_{2}$-balanced and $\mathcal{V}^{\prime}$ is $(1\pm 4\eta)m_{1}$-balanced. Moreover, there is a cycle $C$ on $ts+t=t(s+1)$ vertices such that we can index the family $\mathcal{K}\cup\mathcal{V}^{\prime}$ by $V(C)$, so that $C(\mathcal{V}^{\prime}\cup\mathcal{K})$ is a spanning blow-up of a cycle in $G$, with all clusters of size at least $m_{2}$.
To finish, we subdivide the clusters of $\mathcal{V}^{\prime}$ into $\ell$ clusters of size as close as possible. Since $\mathcal{V}^{\prime}$ is $(1\pm 4\eta)m_{1}$-balanced, the clusters obtained from this subdivision all have size $(1\pm 5\eta)m_{1}/\ell=(1\pm 5\eta)m_{2}$. Together with the clusters of $\mathcal{K}$ we obtain a set family $\mathcal{W}$, which by construction is $(1\pm 5\eta)m_{2}$-balanced. Now, from $C$ we can obtain a new cycle by ‘winding around’ the clusters of $\mathcal{V}^{\prime}$ (using each cycle $C_{i}$ $\ell$ times in each $R_{i}$), whose vertices are in correspondence with $\mathcal{W}$, so that $\mathcal{W}(C)\subseteq G$ is a spanning blow-up of a cycle. This gives the desired $(1\pm 5\eta)m_{2}$-blow-up of a cycle. ∎